OpenStudy (anonymous):

If the length of the longer leg in a 30o-60o-90o triangle is [5\sqrt{3}] ,what is the length of the shorter leg?

OpenStudy (anonymous):

\[\frac{5}{\sqrt{3}} \text{Sin}[30{}^{\circ}]=\frac{5}{2 \sqrt{3}}=1.44338 \]